# Quantum Kernels for Symbolic Sequences Classification

### v1

@@ 2.4 Kernel Functions
A common approach to the solution of the classification problem is to define a similarity measure for pairs of sequences that is consistent with their classes: the sequences belonging to the same class to be closer according to this measure compared to sequences from different classes. A method to formalize the similarity measure is to define a kernel function from the original sequence space to a high dimensional feature space. The intention is the kernel function to linearize the relationships between input and output variables in the feature space. The kernel function approach is a standard which allows the expansion of conventional linear algorithms for classification, regression, and other learning tasks to effectively address nonlinear scenarios.
By contrast, a kernel-based discriminative approach like a support vector classifier starts by properly choosing a kernel, or equivalently a feature map $\phi:X\rightarrow F$, conceptually embeds the data in F and uses the training data to find a linear discriminant in the feature space. Of course, the overall performance will depend strongly on the choice of the kernel: this choice should reflect the prior knowledge about how the data was generated. Lately, there have been many approaches to automatically learn the kernel from data (for instance, [37]). But even in those there is always a prior step where at least a suitable family of kernels must be chosen. Generative kernels try to combine the advantages of both generative and discriminative approaches by introducing generative models on data and use them to devise a kernel. This can be done in many ways: (i) mapping data to points in a probability space and devising a kernel between probability distributions, (ii) considering a fixed probability distribution and study how does it “fit” each data point, (iii) assuming that there is some hidden model that governs the data generation and marginalizing with respect to this model, etc. W
Following the kernel construction approach, we map the input sequences $\mathbf{x}\in X$ to a high-dimensional quantum Hilbert space $H$. The purpose of this mapping, denoted as $\phi:X\rightarrow H$, is to make the output variables linearly separable with respect to their classes. We will use the quantum generator () to define a quantum kernel function $\kappa:X\times X\rightarrow\mathbb{R}$ to compute the inner product in $H$ without the need to explicitly construct its vectors:
$\kappa(x_{1},x_{2})=\langle\phi(x_{1}),\phi(x_{2})\rangle$
The existence of such functions $\kappa$, referred to as kernel functions, is a result of the properties of the inner product structure within a Hilbert space. Mercer’s theorem defines the necessary and sufficient conditions for a function to be a valid kernel function. The kernel functions allow construction of classification and regression learning algorithms that are nonlinear in the input space $X$ while having linear counterparts within the Hilbert space $H$.
2.5 Quantum Hidden Markov Models
@@
  • $\mathcal{H}$ is an $N$-dimensional Hilbert space with space of density operators $D(\mathcal{H})$.
@@
  • $\{T_{a}\}_{a\in\Sigma},\sum_{a\in\Sigma}T_{a}^{\dagger}T_{a}=I_{N}$ is an operator-sum representation of $\mathcal{T}$ in terms of a complete set of Kraus operators.
@@
When the operation $\mathcal{T}=\{T_{a}\}_{a\in\Sigma}$ is applied to a state $\rho\in D(\mathcal{H})$, the model $\mathbf{Q}$ defines measurement outcome $a\in\Sigma$ with probability:
@@
$D^{\mathbf{Q}}_{\rho}=\{P\bigl{[}a|\rho\bigr{]}:a\in\Sigma\}$ (15)
If the operation $\mathcal{T}$ is applied $t$ times starting at the initial state a time series $\mathbf{a}=a_{1}\ldots a_{t}$ will be observed with probability
$P\bigl{[}\mathbf{a}|\rho_{0}\bigr{]}=\operatorname{tr}(T_{\mathbf{a}}\rho_{0}).$ (16)
@@
$\rho_{\mathbf{a}}=\frac{T_{\mathbf{a}}\rho_{0}}{P\bigl{[}\mathbf{a}|\rho_{0}\bigr{]}}.$ (17)
@@
$f^{\mathbf{Q}}(\mathbf{a})=P\bigl{[}\mathbf{a}|\rho_{0}\bigr{]},\forall\mathbf{a}\in\Sigma^{*},$ (18)
@@
$D^{\mathbf{Q}}_{t}=\bigl{\{}{f^{\mathbf{Q}}}(\textbf{a}):\textbf{a}\in{\Sigma}^{t}\bigr{\}}$ (19)
@@
The Definition 1 of a QHMM is based on the concept of a POVM operation acting on a quantum state. The emission of the observable symbols is encoded in the operational elements (Kraus operators) of the quantum operation. This framework provides a convenient way to view QHMMs as channels for quantum information processing. It allows for the analysis of their informational complexity, expressive capacity, and establishes connections to stochastic process languages and the corresponding automata. However, this approach cannot be directly used for implementation of the QHMMs on quantum computing hardware. A critical result in quantum information theory - the Stinespring’s representation theorem  [40] - provides approach to the physical implementation of quantum channels and correspondingly of QHMMs. According to the theorem any quantum channel can be realized as a unitary transformation on a larger system (the combined hidden state and observable systems) followed by a partial trace operation that discards the observable system. An immediate consequence of this result is the following definition of QHMMs in the unitary circuits model of computation:
Definition 2 (Unitary Quantum Hidden Markov Model).
@@
$\mathbf{Q}=\bigl{\{}\Sigma,\mathcal{H}_{S},\mathcal{H}_{E},U,\mathcal{M},R_{0}\bigr{\}}$
@@
The equivalence of the Definition 1 of a QHMM as quantum channel and the unitary representation Definition 2 is defined by the operator-sum representation [39] of the quantum operation $\mathcal{T}$ as follows:
@@ 3 Quantum Generative Kernels
In this section, we introduce measures of similarity for sequences sampled from a stochastic process language. We assume that a generative model of the language (12) is given and define the similarity through this model.
@@ 3.1 Predictive Generative Kernels
The first measure estimates the similarity by evaluating the stochastic distance between the expected future evolutions of the sequences. Since the evolution of a Markovian process depends only of its initial state, it is easy to verify that if two sequences $\mathbf{y}$ and $\mathbf{z}$ define the same quantum sate $\rho$ of the model:
$\rho=\frac{T_{\mathbf{y}}\rho_{0}}{P\bigl{[}{\mathbf{y}|\rho_{0}\bigr{]}}}=\frac{T_{\mathbf{z}}\rho_{0}}{P\bigl{[}{\mathbf{z}|\rho_{0}\bigr{]}}},$
then the sequences $\mathbf{y}$ and $\mathbf{z}$ are equivalent in sense, that they define the same future distributions of the observables.The following proposition generalizes this idea by proving that if two states are close in trace norm, the corresponding forward distributions of the sequences are close in total variation distance.
@@
For any two states $\rho_{1}$ and $\rho_{2}$ of a QHMM $\mathbf{Q}$ the total variation distance of the observable distributions $P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}$ and $P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}$ for $\mathbf{a}\in\Sigma^{l},l>0$ is bounded by the trace distance between the states $\rho_{1}$ and $\rho_{2}$.
@@
The probabilities the same sequence $\mathbf{a}\in\Sigma^{l}$ to be observed at states $\rho_{1}$ and $\rho_{2}$ respectively are (16)
$P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}=\operatorname{tr}(T_{\mathbf{a}}\rho_{1})$ (23)
@@
$P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}=\operatorname{tr}(T_{\mathbf{a}}\rho_{2})$ (24)
@@
$P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}-P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}=\operatorname{tr}(T_{\mathbf{a}}(\rho_{1}-\rho_{2}))$ (25)
Since the operation $T_{a}$ is trace non-increasing we have:
$P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}-P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}\leq\operatorname{tr}(\rho_{1}-\rho_{2}).$ (26)
If we assume that $P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}\geq P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}$ then
$\bigl{|}P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}-P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}\bigr{|}\leq 2\mathcal{D}(\rho_{1},\rho_{2}).$ (27)
The total variation distance between the distributions $P_{1}(\mathbf{a})=P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}$ and $P_{2}(\mathbf{a})=P\bigl{[}\mathbf{a}|\rho_{2}\bigr{]}$ is
$\delta(P_{1},P_{2})=\sup_{\mathbf{a}\in\Sigma^{l}}\bigl{|}P_{1}(\mathbf{a})-P_{2}(\mathbf{a})\bigr{|}$ (28)
@@
This proposition implies that when, at a specific time $t=t_{0}$, two hidden states of the model are close in trace norm, the distributions of the observable processes originated from these states will be close in total variation distance for $t>t_{0}$.
To understand the importance of this fact in the context of kernels for discriminative tasks, let’s consider a specific classification problem. In this problem, each observed sequence $\mathbf{y}\in L^{Q}$ is classified based on its $k$-step future evolution. More formally, if at time $t_{0}$ we have observed the sequence $\mathbf{y}$ and at time $t_{0}+k$ we have observed the sequence $\mathbf{ya}$, then the class of $\mathbf{y}$ depends only of some feature of the sequence $\mathbf{a}$. We assume that the class can be computed by a function $f:\Sigma^{k}\rightarrow C$. As the sequence $\mathbf{y}$ is sampled from the stochastic language $L^{Q}$ described by the generative model $Q$ (12), the mapping function (11) can be defined as follows :
$p(\mathbf{y},c)=\sum_{\mathbf{a}\in\Sigma^{k}}P\bigl{[}\mathbf{a}|\mathbf{y}\bigr{]}\cdot I_{f({\mathbf{a})=c}},$ (30)
where
$P\bigl{[}\mathbf{a}|\mathbf{y}\bigr{]}=P\bigl{[}\mathbf{a}|\rho_{\mathbf{y}}\bigr{]},$
and from (14)
$\rho_{\mathbf{y}}=\frac{T_{\mathbf{y}}\rho_{0}}{P\bigl{[}{\mathbf{y}|\rho_{0}\bigr{]}}},$
The mapping function (30) defines a distribution over the class labels:
$CP_{\mathbf{y}}(c)=\bigl{\{}p(\mathbf{y},c):c\in C\bigr{\}}$
Any other sequence $\mathbf{z}\in L^{Q}$ defines distribution $CP_{\mathbf{z}}$ at total variation distance w.r.t $CP_{\mathbf{y}}$:
$\delta(CP_{\mathbf{y}},CP_{\mathbf{z}})=\sup_{c\in C}\bigl{|}CP_{\mathbf{y}}(c)-CP_{\mathbf{z}}(c)\bigr{|}$ (31)
$\delta(CP_{\mathbf{y}},CP_{\mathbf{z}})=\sup_{c\in C}\sum_{\mathbf{a}\in\Sigma^{k}}\bigl{|}P\bigl{[}\mathbf{a}|\mathbf{y}\bigr{]}-P\bigl{[}\mathbf{a}|\mathbf{z}\bigr{]}\bigr{|}\cdot I_{f({\mathbf{a})=c}}$ (32)
From Proposition 1:
$\delta(CP_{\mathbf{y}},CP_{\mathbf{z}})\leq 2\mathcal{D}(\rho_{\mathbf{y}},\rho_{\mathbf{z}})\cdot\sup_{c\in C}\sum_{\mathbf{a}\in\Sigma^{k}}I_{f({\mathbf{a})=c}}$ (33)
$\delta(CP_{\mathbf{y}},CP_{\mathbf{z}})\leq C\cdot\mathcal{D}(\rho_{\mathbf{y}},\rho_{\mathbf{z}}),$ (34)
where
$C=2\sup_{c\in C}\sum_{\mathbf{a}\in\Sigma^{k}}I_{f(\mathbf{a})=c}$
is a constant depending on the structure of the classification task.
This result demonstrates that the variation distance between the classification distributions of two sequences is bounded by the trace distance between the corresponding quantum states. This motivates us to define the similarity between every pair of time series $\mathbf{t}_{1}$ and $\mathbf{t}_{2}$ by mapping them to the quantum states $\rho_{1}$ and $\rho_{2}$ of the model (1), and calculating the trace distance between these states. It is important to emphasize that this measure of similarity pertains to the future evolution of the time series. We define the predictive generative kernel as follows:
$\kappa(\mathbf{y},\mathbf{z})=\mathcal{D}\bigl{(}\phi(\mathbf{y}),\phi(\mathbf{z})\bigr{)},\mathbf{y},\mathbf{z}\in L^{Q}$ (35)
where the mapping $\phi(\mathbf{y})$ of a sequence $\mathbf{y}$ to quantum state $\rho_{\mathbf{y}}$ is defined as:
@@
Since the Bures metric is symmetric and positive semi-definite it defines a valid kernel:
$\kappa(\mathbf{y},\mathbf{z})=\mathcal{B}\bigl{(}\phi(\mathbf{y}),\phi(\mathbf{z})\bigr{)},\mathbf{y},\mathbf{z}\in L^{Q}.$ (36)

### v2

@@ 2.4 Kernel Functions
A common approach to solving classification problems is to establish a similarity measure for pairs of examples that is consistent with their respective classes: examples within the same class should be considered closer according to this measure compared to examples from different classes. The similarity measure can be formalized as a kernel function, which computes the similarity or distance between pairs of examples in the input domain. We will assume that the example domain is a stochastic process language $L$ as discussed in Section 2.1. A kernel function is formally defined as follows:
$K:L\times L\rightarrow\mathbb{R},$ (13)
where $K$ is a symmetric, positive semi-definite function.
The idea behind this definition is to map the examples into a high-dimensional feature space where the data becomes more separable or even linearly separable, and then to compute similarity within that space. Let’s consider an implicit or explicit map of the input examples to a high-dimensional Hilbert space $H$:
$\phi:L\rightarrow H.$ (14)
If $\kappa$ is a symmetric positive semi-definite measure in the Hilbert space:
$\kappa:H\times H\rightarrow\mathbb{R},$ (15)
then the kernel function is specified as:
$K(x_{1},x_{2})=\kappa(\phi(x_{1}),\phi(x_{2}))$ (16)
We construct a kernel function using the state-space of the stochastic generative model (6). The model assigns probability to every observable sequence by equation (8). The model also defines a mapping between the examples $\mathbf{y}=y_{1}\ldots y_{t}$ and the vectors of probabilities of observing an example $\mathbf{y}$ from each of the n possible states or inference paths:
$\phi_{\mathbf{M}}(\textbf{y})=T_{y_{t}}\ldots T_{y_{1}}\textbf{x}_{0}=T_{\textbf{y}}\textbf{x}_{0}$ (17)
construction of classification and regression learning algorithms that are nonlinear in the input space $L$ while having linear counterparts within the Hilbert space $H$.
In the next session we assume the existence of a quantum generative model for the language $L$ and apply this approach in the Hilbert space of the model.
3 Quantum Hidden Markov Models
@@
  • $\mathcal{H}$ is an $N$-dimensional Hilbert space. It defines the state-space of the model $\mathbf{Q}$ as the set of associated density operators $D(\mathcal{H})$.
@@
  • $\{T_{a}\}_{a\in\Sigma},\sum_{a\in\Sigma}T_{a}^{\dagger}T_{a}=I_{N}$ is an operator-sum representation of $\mathcal{T}$ in terms of a complete set of Kraus operators $T_{a}$.
@@
The space $D(\mathcal{H})$ is a convex set, representing statistical ensembles of quantum states, which consists of all positive semi-definite operators $\rho$ on $\mathcal{H}$ with $\operatorname{tr}(\rho)=1$. When the map $\mathcal{T}=\{T_{a}\}_{a\in\Sigma}$ is applied to a state $\rho\in D(\mathcal{H})$, the model $\mathbf{Q}$ defines measurement outcome $a\in\Sigma$ with probability:
@@
$D^{\mathbf{Q}}_{\rho}=\{P\bigl{[}a|\rho\bigr{]}:a\in\Sigma\}.$ (21)
If the operation $\mathcal{T}$ is applied $t$ times starting at the initial state a time series $\mathbf{y}=y_{1}\ldots y_{t}$ will be observed with probability
$P\bigl{[}\mathbf{y}|\rho_{0}\bigr{]}=\operatorname{tr}(T_{\mathbf{y}}\rho_{0}).$ (22)
@@
$\rho_{\mathbf{y}}=\frac{T_{\mathbf{y}}\rho_{0}}{P\bigl{[}\mathbf{y}|\rho_{0}\bigr{]}},$ (23)
@@
$f^{\mathbf{Q}}(\mathbf{y})=P\bigl{[}\mathbf{y}|\rho_{0}\bigr{]},\forall\mathbf{y}\in\Sigma^{*}.$ (24)
@@
$D^{\mathbf{Q}}_{t}=\bigl{\{}{f^{\mathbf{Q}}}(\textbf{y}):\textbf{y}\in{\Sigma}^{t}\bigr{\}}$ (25)
@@
The Definition 1 of a QHMM is based on the concept of a POVM operation acting on a quantum state. The emission of the observable symbols is encoded in the operational elements (Kraus operators) of the quantum operation. This framework provides a convenient way to view QHMMs as channels for quantum information processing. It allows for the analysis of their informational complexity, expressive capacity, and establishes connections to stochastic process languages and the corresponding automata. However, this approach cannot be directly used for implementation of the QHMMs on quantum computing hardware. A critical result in quantum information theory — the Stinespring’s representation theorem  [51] — provides approach to the physical implementation of quantum channels and correspondingly of QHMMs. According to the theorem any quantum channel can be realized as a unitary transformation on a larger system (the combined hidden state and observable systems) followed by a partial trace operation that discards the observable system. An immediate consequence of this result is the following definition of QHMMs in the unitary circuits model of computation:
Definition 2 (Unitary quantum hidden Markov model [37]).
@@
$\mathbf{Q}=\bigl{\{}\Sigma,\mathcal{H}_{S},\mathcal{H}_{E},U,\mathcal{M},R_{0}\bigr{\}}$ (26)
@@
[…]
@@
{quantikz}
[…]
Figure 1: Implementation of QHMM.
$State$ Direction Bias 00 1 2 3 $P[0|State]$ $P[1|State]$
0 Bear - Tendency Down 0.50 0.10 0.15 0.25 0.80 0.20
1 Bull - Tendency Up 0.10 0.50 0.25 0.15 0.20 0.80
2 Transition to Bear 0.25 0.15 0.50 0.10 0.40 0.60
3 Transition to Bull 0.15 0.25 0.10 0.50 0.60 0.40
Table 1: Hidden states descriptions, transition probabilities, and observation probabilities.
Figure 2: Left: Market hidden states transition graph. Right: Distributions of observed sequences.
Figure 3: QHMM defining market distribution
Figure 4: Targeted and Learned Market Process Language Distributions
Example 1.
In [37] we discussed a stochastic process language generated by a simple classic hidden Markov model (HMM) of directional market price movements. The market is assumed to have four hidden states and observable symbols ${0,1}$ corresponding to price move down and up. Table 1 provides the transition and emission probabilities of the model. The hidden states transition graph and the distributions defining the stochastic process language are presented on Figure 2. In [37] we proved that every classical HMM of $n$ states can be simulated by a QHMM (26) in Hilbert space with dimension $\sqrt{n}$. Following this result we identified a generative model — a QHMM with one state qubit and one emission qubit (Figure 3) — which reproduces exactly the distributions of the market process language as shown on Figure 4. The first part of the circuit prepares a maximally mixed initial state, the middle part performs the hidden state transition, and at the end is the measurement of the emission subsystem in a learned orthonormal basis.
@@ 4 Quantum Generative Kernels
In this section, we introduce sequence similarity measures based on the quantum generative model defined by equations (18) and (25). We assume that a QHMM which defines the distribution set (25) of the example space is specified either as a quantum channel (18) or unitary model (26).
@@ 4.1 Predictive Generative Kernels
The predictive similarity measure is defined by the stochastic distance between the expected future evolutions of the sequences. Since the evolution of a Markovian process depends only on its current state, it is easy to verify that if two sequences $\mathbf{y^{1}}$ and $\mathbf{y^{2}}$ define the same quantum sate $\rho$ of the model (23):
$\rho=\frac{T_{\mathbf{y^{1}}}\rho_{0}}{P\bigl{[}{\mathbf{y^{1}}|\rho_{0}\bigr{]}}}=\frac{T_{\mathbf{y^{2}}}\rho_{0}}{P\bigl{[}{\mathbf{y^{2}}|\rho_{0}\bigr{]}}},$
then the sequences $\mathbf{y^{1}}$ and $\mathbf{y^{2}}$ are considered equivalent in the sense, that they define the same future distributions of the observables.
The following proposition generalizes this idea by proving that if two states are close in trace norm, the corresponding forward distributions of the sequences are close in total variation distance.
@@
For any two states $\rho_{1}$ and $\rho_{2}$ of a QHMM $\mathbf{Q}$ the total variation distance of the observable distributions $P\bigl{[}\mathbf{z}|\rho_{1}\bigr{]}$ and $P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}$ for $\mathbf{z}\in\Sigma^{k},k>0$ is bounded by the trace distance between the states $\rho_{1}$ and $\rho_{2}$.
@@
The probabilities the same sequence $\mathbf{z}\in\Sigma^{k}$ to be observed at states $\rho_{1}$ and $\rho_{2}$ respectively are (22)
$P\bigl{[}\mathbf{z}|\rho_{1}\bigr{]}=\operatorname{tr}(T_{\mathbf{z}}\rho_{1}),$ (30)
@@
$P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}=\operatorname{tr}(T_{\mathbf{z}}\rho_{2}).$ (31)
@@
$P\bigl{[}\mathbf{z}|\rho_{1}\bigr{]}-P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}=\operatorname{tr}(T_{\mathbf{z}}(\rho_{1}-\rho_{2})).$ (32)
Since the operation $T_{z}$ is trace non-increasing we have:
$P\bigl{[}\mathbf{z}|\rho_{1}\bigr{]}-P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}\leq\operatorname{tr}(\rho_{1}-\rho_{2}).$ (33)
If we assume that $P\bigl{[}\mathbf{z}|\rho_{1}\bigr{]}\geq P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}$ then
$\bigl{|}P\bigl{[}\mathbf{z}|\rho_{1}\bigr{]}-P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}\bigr{|}\leq 2\mathcal{D}(\rho_{1},\rho_{2}).$ (34)
The total variation distance between the distributions $P_{1}(\mathbf{z})=P\bigl{[}\mathbf{a}|\rho_{1}\bigr{]}$ and $P_{2}(\mathbf{z})=P\bigl{[}\mathbf{z}|\rho_{2}\bigr{]}$ is
$\delta(P_{1},P_{2})=\sup_{\mathbf{z}\in\Sigma^{k}}\bigl{|}P_{1}(\mathbf{z})-P_{2}(\mathbf{z})\bigr{|}$ (35)
Since (34) is valid for any $\mathbf{z}\in\Sigma^{k}$, then from (35) follows:
@@
This proposition motivates us to define the similarity between every sequence pair $\mathbf{y}^{1}$ and $\mathbf{y}^{2}$ by mapping them to the quantum states $\rho_{1}$ and $\rho_{2}$ of the model (18), and calculating the trace distance between these states. Further in this chapter we will proof that this measure of similarity pertains to the future evolution of the time series.
We define the predictive generative kernel as follows:
Definition 3 (Predictive generative quantum kernel).
For every sequences $\mathbf{y}^{1}$ and $\mathbf{y}^{2}$ in the example domain $L^{\mathbf{Q}}$ of a QHMM $\mathbf{Q}$ a predictive kernel $\kappa_{p}:L^{\mathbf{Q}}\times L^{\mathbf{Q}}\rightarrow\mathbb{R}$ is the function:
$\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})=\mathcal{D}\bigl{(}\phi(\mathbf{y}^{1}),\phi(\mathbf{y}^{2})\bigr{)},\mathbf{y}^{1},\mathbf{y}^{2}\in L^{Q}$ (37)
where the mapping $\phi(\mathbf{y})$ of a sequence $\mathbf{y}$ to quantum state $\rho_{\mathbf{y}}$ is defined as (23) :
@@
Critical feature of the kernel is that it reflects the similarity of sequences with respect of their probabilistic classification maps (12) as it is demonstrated by the following Proposition.
Proposition 2.
For any two sequences $\mathbf{y}^{1}$ and $\mathbf{y}^{2}$ the variation distance of the classification maps (12) $p_{p}(\mathbf{y}^{1},k,c)$ and $p_{p}(\mathbf{y}^{2},k,c),k>0$ is bounded by the predictive kernel distance $\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2})$.
Proof.
The variation distance of the class distributions (12) for the sequences is defined as
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=\sup_{c\in C}\bigl{|}p_{p}(\mathbf{y}^{1},k,c)-p_{p}(\mathbf{y}^{2},k,c)\bigr{|}$ (38)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=\sup_{c\in C}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{1}]P[c|\mathbf{z}])-\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{2}]P[c|\mathbf{z}])\bigr{|}$ (39)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=\sup_{c\in C}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}P[c|\mathbf{z}](P[\mathbf{z}|\mathbf{y}^{1}])-P[\mathbf{z}|\mathbf{y}^{2}])\bigr{|}$ (40)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))=c^{*}\bigl{|}\sum_{\mathbf{z}\in\Sigma^{k}}(P[\mathbf{z}|\mathbf{y}^{1}])-P[\mathbf{z}|\mathbf{y}^{2}])\bigr{|}$ (41)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))\leq c^{*}\sum_{\mathbf{z}\in\Sigma^{k}}(\bigl{|}P[\mathbf{z}|\mathbf{y}^{1}])-P[\mathbf{z}|\mathbf{y}^{2}]\bigr{|})$ (42)
From Proposition 1
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))\leq c^{*}|\Sigma^{k}|2\mathcal{D}(\rho_{1},\rho_{2})$ (43)
$\delta(p_{p}(\mathbf{y}^{1},k,c),p_{p}(\mathbf{y}^{2},k,c))\leq 2c^{*}|\Sigma^{k}|\kappa_{p}(\mathbf{y}^{1},\mathbf{y}^{2}),$ (44)
where $c^{*}\leq 1$ is a constant dependent on the particular probabilistic classification map $P[c|\mathbf{z}]$.
@@
Since the Bures metric is symmetric and positive semi-definite the following equality defines a valid kernel:
$\kappa_{p}(\mathbf{y},\mathbf{z})=\mathcal{B}\bigl{(}\phi(\mathbf{y}),\phi(\mathbf{z})\bigr{)},\mathbf{y},\mathbf{z}\in L^{Q}.$ (45)